OpenStudy (anonymous):

what is the point of intersection? 10y=x^2 x^2-6=-2y

terenzreignz (terenzreignz):

I suggest isolating y on both equations.

OpenStudy (anonymous):

I would try elimination.

terenzreignz (terenzreignz):

So that you have \[\huge y=\frac{x^2}{10}\]and\[\huge y=3-\frac{x^2}{2}\]

terenzreignz (terenzreignz):

And make these two functions of x equal to each other, and solve. Good luck :)

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

Elimination gives you \[ x^2-6-(x^2)=-2y-(10y) \\ -6=-12y \]

OpenStudy (anonymous):

y=1/2 there is an intersection right?

terenzreignz (terenzreignz):

Try solving this \[\huge \frac{x^2}{10}=3-\frac{x^2}{2}\]

terenzreignz (terenzreignz):

\[\huge x^2 = 30 - 5x^2\]